# From System 1 to System 2: A Survey of Reasoning Large Language Models

### v2

@@ 3.3 Evolutionary of Reasoning LLMs
Agent WebShop [254] Open-End NeurIPS 2022 English 1,600 Hybrid
@@
In the early stages, reasoning LLMs primarily focused on enhancing pre-trained LLMs with external reasoning algorithms, without altering the underlying model parameters. Approaches such as Tree of Thoughts [265] and Reasoning via Planning [14] utilized LLMs-driven Breadth-First Search, Depth-First Search, and MCTS [266, 98, 128, 125] to simulate human-like reasoning processes. These methods represented reasoning as tree or graph traversals, where intermediate reasoning states were depicted as nodes, and various reasoning strategies produced distinct reasoning paths. The final decision was made through additional voting mechanisms [3] or Monte Carlo-based value estimation to identify the optimal path.
@@
With the introduction of models like o1 [29] and QwQ [118], external reasoning paradigms were internalized within the LLM’s context. These models initially performed exploratory macro-planning to generate an initial reasoning path, followed by contextual exploration of alternative paths. Through mechanisms like “Rethink” and “Verification”, these models produced extended reasoning chains. To replicate this internalized capability, STILL-1 [266] linearized tree search outputs into long reasoning chains with attributes such as “Rethink”, “Wait”, and “Explore New Path”. Similarly, STILL-2 [53] and sky-T1 [119] synthesized long reasoning chains using distillation techniques. However, the linearized reasoning chains derived from search-based methods struggled to match the quality of those produced by distillation approaches.
Recent advancements, including DeepSeek-R1 [31] and Kimi-k1.5 [215], have demonstrated the potential of RL to enhance models like DeepSeek-V3 [17], resulting in the emergence of complex behaviors such as long reasoning chains, reflective reasoning, and advanced planning capabilities. Remarkably, these sophisticated behaviors were achieved through simple RL scaling. SimpleRL [123] sought to replicate these capabilities using a streamlined pipeline and minimal codebase, while R1V [234] explored the development of multimodal reasoning models based on multimodal foundation architectures.
@@ 4.1 Benchmark Categories
We categorize reasoning benchmarks by task type, which can be broadly divided into math, code, scientific, agent, medical, and multimodal reasoning. The detailed statistics for these benchmarks are presented in Table VI.
@@ 4.1.1 Benchmark Introduction
    Multimodal Reasoning: Multimodal reasoning, such as benchmarks of MMMU [260] and MathVista [261], requires cross-modal thinking in combination with text and images. Especially for those visual-centered problems, in benchmarks MathVision [262], MathVerse [269], CMMaTH [263], and PGPS9K [264], put forward higher requirements for reasoning LLMs.
@@ 4.2 Evaluation Metrics
TABLE VII: Performance of Different Models, including Basic LLMs and Reasoning LLMs, on Plain Text Benchmarks. The red denotes the highest result, and the blue denotes the second highest result.
@@
Basic LLMs GPT-4o [16] 9.3 74.6 34.2 23.6 38.8 87.2 49.9
@@ 4.3.1 Performance on Plain Text Benchmarks
As shown in Table VII, reasoning LLMs, such as DeepSeek-R1 [31] and OpenAI-o1/o3 [29, 30], demonstrate exceptional performance across a wide range of tasks, including math, coding, and other general tasks. These models achieve high scores on multiple plain-text benchmarks, such as AIME 2024, MATH-500, and LiveCodeBench, showcasing their robust text-based reasoning abilities. In contrast, foundational LLMs, like GPT-4o [62], Claude-3.5-Sonnet [272], and DeepSeek-V3 [17], generally perform less effectively than reasoning LLMs, particularly in math and coding tasks (e.g., AIME 2024 and Codeforces). For example, OpenAI-o1 outperforms GPT-4o by 69.9% and 73% on these tasks, respectively. Moreover, DeepSeek-R1, based on the DeepSeek-V3 architecture, surpasses its predecessor on all benchmarks, further highlighting the advantages of the reasoning LLMs.
@@ 4.3.2 Performance on Multimodal Benchmarks
As shown in Table VIII, reasoning LLMs continue to excel in multimodal tasks. OpenAI-o1 [29] performs strongly in vision tasks, achieving the highest score of 77.3% on MMMU and outperforming its corresponding foundational LLM, GPT-4o [62], by 7.2% on MathVista. However, the performance improvement in multimodal tasks is less pronounced compared to text-only tasks. This can be attributed in part to the limitations of current multimodal reasoning LLM techniques, as well as the lack of sufficient datasets to fully assess the multimodal capabilities of reasoning LLMs.
TABLE VIII: Performance of Models, including Basic LLMs and Reasoning LLMs, on Multimodal Benchmarks. The red denotes the highest result, and the blue denotes the second highest result.
@@
Basic LLMs GPT-4o [16] 69.1 63.8 30.4 25.9
Claude-3.5-Sonnet [272] 70.4 65.3 35.6 -
@@
Reasoning LLMs LLaVA-CoT [197] - 54.8 - -
@@
Kimi k1.5 long-CoT [215] 70.0 74.9 - -
OpenAI-o1 [29] 77.3 71.0 - -

### v3

@@ 3.3 Evolutionary of Reasoning LLMs
Putnam-AXIOM [299] Open-End NeurIPS 2024 English 236 Competition
PRM800K [178] Open-End ICLR 2024 English 800,000 Hybrid
FrontierMath [300] Open-End ArXiv 2024 English - Expert
ProcessBench [301] Open-End ArXiv 2024 English 3400 Competition
LiveBench [302] Open-End ICLR 2025 English Frequently Updated Expert
AIME 2025 Open-End Hugging Face English 13 Competition
ThinkBench [303] Hybrid ArXiv 2025 English 2,912 Expert
MATH-Perturb [304] Open-End ArXiv 2025 English 279 Competition
ZebraLogic [305] Open-End ArXiv 2025 English 1,000 Hybrid
QuestBench [306] Choice ArXiv 2025 English 38,882 Hybrid
Math-RoB [307] Open-End ArXiv 2025 English - High School
GSM-Ranges [308] Open-End ArXiv 2025 English - Middle School
@@
MHPP [316] Open-End ArXiv 2024 English 210 Expert
RewardBench [317] Hybrid ArXiv 2024 English - Hybrid
MR-Ben [318] Open-End NeurIPS 2024 English 5,975 Hybrid
ReaLMistake [319] Open-End COLM 2024 English - Expert
CriticBench [320] Open-End ACL 2024 English - Hybrid
JudgeBench [321] Open-End ArXiv 2024 English - Hybrid
TPBench [322] Open-End ArXiv 2025 English 57 University
ProBench [323] Open-End ArXiv 2025 English/Chinese 790 Competition
EquiBench [324] Open-End ArXiv 2025 English 2,400 Hybrid
SuperGPQA [325] Choice ArXiv 2025 English 26,529 University
Sys2Bench [326] Open-End ArXiv 2025 English - Hybrid
PRMBench [327] Open-End ArXiv 2025 English 6,216 Expert
DeltaBench [328] Open-End ArXiv 2025 English - Expert
FINEREASON [329] Open-End ArXiv 2025 English - Expert
Agent ARC [330] Open-End ArXiv 2019 Symbolic 1,000 Expert
WebShop [331] Open-End NeurIPS 2022 English 1,600 Hybrid
SciWorld [332] Open-End EMNLP 2022 English 7,200 Hybrid
@@
Osworld [335] Open-End NeurIPS 2024 English 369 Hybrid
GAMABench [336] Open-End ArXiv 2024 English - Hybrid
Mle-bench [337] Open-End ArXiv 2025 English - Competition
ToolComp [338] Open-End ArXiv 2025 English - Hybrid
Mobile-Agent-E [339] Open-End ArXiv 2025 English 25 Hybrid
Text2World [340] Open-End ArXiv 2025 English - Hybrid
WebGames [341] Open-End ArXiv 2025 English 50 Hybrid
Ui-r1 [342] Open-End ArXiv 2025 English 136 Hybrid
@@
MEDEC [345] Open-End ArXiv 2024 English 3,848 Expert
MedXpertQA [346] Choice ArXiv 2025 English 4,460 Expert
TABLE VII: Statistics of benchmarks for reasoning MLLMs.
Domain Benchmark Question Type Venue Language Size Level
@@
ZeroBench [352] Open-End ArXiv 2025 English 100/334 Impossible
MME-CoT [353] Hybrid ArXiv 2025 English 1,130 Hybrid
MM-IQ [354] Choice ArXiv 2025 English/Chinese 2,710 Hybrid
Multimodal RewardBench [355] Hybrid ArXiv 2025 English 5,211 Hybrid
GRAB [356] Open-End ArXiv 2024 English 2,710 Hybrid
SciFIBench [357] Choice ArXiv 2024 English 2,000 Challenging
MV-MATH [358] Hybrid ArXiv 2025 English 2,009 Middle/High School
ScienceQA [359] Choice NeurIPS 2022 English 21,000 Hybrid
Plot2Code [360] Open-End ArXiv 2024 English 132 Hybrid
M3CoT [361] Open-End ArXiv 2024 English - Hybrid
PUZZLEVQA [362] Hybrid ACL 2024 English - Hybrid
MolPuzzle [363] Hybrid NeurIPS 2024 English 23,000 Expert
HumanEval-V [364] Open-End ArXiv 2024 English 108 Hybrid
CoMT [365] Choice AAAI 2025 English 3,853 Hybrid
ChartMimic [366] Open-End ArXiv 2024 English 4,800 Expert
OlympicArena [367] Hybrid NeurIPS 2024 English/Chinese 11,163 Competition
CVQA&CPVQA [368] Open-End ArXiv 2025 English 1664 Hybrid
ENIGMAEVAL [369] Open-End ArXiv 2025 English 1184 Challenging
CODE-VISION [370] Open-End ArXiv 2025 English 438 Competition
MKRC [371] Open-End ArXiv 2025 English 7010 Hybrid
MMSciBench [372] Hybrid ArXiv 2025 English/Chinese 4,482 High School
LEGO-Puzzles [373] Hybrid ArXiv 2025 English 1,100 Expert
JustLogic [374] Open-End ArXiv 2025 English 7,000 Expert
HUMANITY’S LAST EXAM [375] Hybrid ArXiv 2025 English 2,700 Challenging
DivIL [376] Hybrid TMLR English - Hybrid
ErrorRadar [377] Open-End ArXiv 2024 English 2,500 Middle/High School
In the early stages, reasoning LLMs primarily focused on enhancing pre-trained LLMs with external reasoning algorithms, without altering the underlying model parameters. Approaches such as Tree of Thoughts [378, 379] and Reasoning via Planning [14] utilized LLMs-driven Breadth-First Search, Depth-First Search, and MCTS [380, 98, 129, 127] to simulate human-like reasoning processes. These methods represented reasoning as tree or graph traversals, where intermediate reasoning states were depicted as nodes, and various reasoning strategies produced distinct reasoning paths. The final decision was made through additional voting mechanisms [3] or Monte Carlo-based value estimation to identify the optimal path.
@@
With the introduction of models like o1 [29] and QwQ [120], external reasoning paradigms were internalized within the LLM’s context. These models initially performed exploratory macro-planning to generate an initial reasoning path, followed by contextual exploration of alternative paths. Through mechanisms like “Rethink” and “Verification”, these models produced extended reasoning chains. To replicate this internalized capability, STILL-1 [380] linearized tree search outputs into long reasoning chains with attributes such as “Rethink”, “Wait”, and “Explore New Path”. Similarly, STILL-2 [53] and sky-T1 [121] synthesized long reasoning chains using distillation techniques. Approaches like Virgo [382] have attempted to distill text-based slow-thinking reasoning into multimodal LLMs; their performance improvements in tasks such as MathVision [349], which demand detailed visual understanding, have been marginal. However, the linearized reasoning chains derived from search-based methods struggled to match the quality of those produced by distillation approaches. Additionally, extending slow-thinking reasoning capabilities from text-based domains to multimodal contexts remains a significant challenge, especially in tasks requiring fine-grained perception [117, 383, 384, 385].
Recent advancements, including DeepSeek-R1 [31] and Kimi-k1.5 [255], have demonstrated the potential of RL to enhance models like DeepSeek-V3 [17], resulting in the emergence of complex behaviors such as long reasoning chains, reflective reasoning, and advanced planning capabilities. Remarkably, these sophisticated behaviors were achieved through simple RL scaling. SimpleRL [125] sought to replicate these capabilities using a streamlined pipeline and minimal codebase, while R1V [274] explored the development of multimodal reasoning models based on multimodal foundation architectures. However, though R1 has been proven to significantly enhance reasoning abilities in the LLMs field, there remain many challenges to explore in the MLLMs domain. These include difficulties in maintaining a consistent slow thinking process when handling complex visual inputs and achieving training benefits comparable to those acquired through unimodal RL.
@@ 4.1 Benchmark Categories
We categorize reasoning benchmarks by task type, which can be broadly divided into math, code, scientific, agent [386], medical, and multimodal reasoning. The detailed statistics for these benchmarks are presented in Table VII.
@@ 4.1.1 Benchmark Introduction
    Multimodal Reasoning: Multimodal reasoning, such as benchmarks of MMMU [347] , MathVista [348], MME-CoT [353] and MM-IQ [354], requires cross-modal thinking in combination with text and images. Especially for those visual-centered problems, in benchmarks MathVision [349], MathVerse [388], CMMaTH [350] , PGPS9K [351], SciFIBench [357] and GRAB [356], put forward higher requirements for reasoning MMs.
@@ 4.2 Evaluation Metrics
TABLE VIII: Performance of Different Models, including Basic LLMs and Reasoning LLMs, on Plain Text Benchmarks. The red denotes the highest result, and the blue denotes the second highest result.
@@
Basic LLMs Gemini-2.5-Pro [390] 92.0 - 70.4 - 63.8 81.7 84.0
Gemini-2.0-Pro [391] - 91.8 36.0 - - 86.5 64.7
GPT-4o [16] 9.3 74.6 34.2 23.6 38.8 87.2 49.9
@@
OpenAI o3-mini [30] 87.3 97.9 84.6 - 49.3 86.9 79.7
Seed-Thinking-v1.5 [398] 86.7 - 64.9 55.0 47.0 87.0 77.3
DeepSeek-R1-Distill-Qwen-1.5B [399] 28.9 83.9 16.9 - - - 33.8
DeepSeek-R1-Distill-Qwen-7B[399] 55.5 92.8 37.6 - - - 49.1
DeepSeek-R1-Distill-Qwen-14B [399] 69.7 93.9 53.1 - - - 59.1
DeepSeek-R1-Distill-Qwen-32B[399] 72.6 94.3 57.2 - - - 62.1
DeepSeek-R1-Distill-Llama-8B [399] 50.4 89.1 39.6 - - - 49.0
DeepSeek-R1-Distill-Llama-70B [399] 70.0 94.5 57.5 - - - 65.2
@@ 4.3.1 Performance on Plain Text Benchmarks
As shown in Table VIII, reasoning LLMs, such as DeepSeek-R1 [31] and OpenAI-o1/o3 [29, 30], demonstrate exceptional performance across a wide range of tasks, including math, coding, and other general tasks. These models achieve high scores on multiple plain-text benchmarks, such as AIME 2024, MATH-500, and LiveCodeBench, showcasing their robust text-based reasoning abilities. In contrast, foundational LLMs, like GPT-4o [62], Claude-3.5-Sonnet [392], and DeepSeek-V3 [17], generally perform less effectively than reasoning LLMs, particularly in math and coding tasks (e.g., AIME 2024 and Codeforces). For example, OpenAI-o1 outperforms GPT-4o by 69.9% and 73% on these tasks, respectively. Moreover, DeepSeek-R1, based on the DeepSeek-V3 architecture, surpasses its predecessor on all benchmarks, further highlighting the advantages of the reasoning LLMs.
@@ 4.3.2 Performance on Multimodal Benchmarks
As shown in Table IX, reasoning LLMs continue to excel in multimodal tasks. OpenAI-o1 [29] performs strongly in vision tasks, achieving the highest score of 78.2% on MMMU and outperforming its corresponding foundational LLM, GPT-4o [62], by 7.2% on MathVista. However, the performance improvement in multimodal tasks is less pronounced compared to text-only tasks. This can be attributed in part to the limitations of current multimodal reasoning LLM techniques, as well as the lack of sufficient datasets to fully assess the multimodal capabilities of reasoning LLMs.
TABLE IX: Performance of Models, including Basic LLMs and Reasoning LLMs, on Multimodal Benchmarks. The red denotes the highest result, and the blue denotes the second highest result.
@@
Basic LLMs Claude-3.5-Sonnet [392] 68.3 65.3 38.0 -
GPT-4o [16] 69.1 63.8 30.4 25.9
@@
Llama 4 Maverick [402] 73.4 73.7 - -
Claude-3.7-Sonnet [393] 75.0 - - -
Gemini 2.5 Pro[390] 81.7 - - -
Reasoning LLMs Skywork-R1V-38B [282] 69.0 67.5 - -
LLaVA-CoT [227] - 54.8 - -
Kimi k1.5 long-CoT [255] 70.0 74.9 - -
Qwen2.5-VL-72B [403] 70.2 74.8 38.1 -
@@
Kimi k1.6 preview - 80.0 53.3 -
MMR1-Math-v0-7B [246] - 71.0 30.2 -
MM-EUREKA [247] - 64.2 26.6 37.3
R1-Onevision-7B [253] - 64.1 29.9 -
Doubao-1.5-pro [405] 73.8 78.8 48.6 48.5
OpenAI-o1 [29] 78.2 71.0 - -
OpenAI o4-mini [406] 81.6 84.3 - -
@@ 4.3.3 Summary
5 Extended Techniques
5.1 Advanced Infrastructure
In the realm of System-2 paradigms, reinforcement learning stands as a pivotal methodology to elevate the performance of reasoning models, finding ubiquitous employment in LLM, MLLM, and agentic applications such as Agent-R1 [407] and UI-R1 [408]. Recent advancements have seen state-of-the-art RL frameworks like VeRL [409], OpenRLHF [410], and ReaLHF [411] emerge as dominant tools in this domain. Despite their divergent architectural designs, these frameworks adhere to a canonical RLHF workflow comprising three core stages:
  • Generative Rollout Phase: The actor network generates outputs in response to a batch of input prompts. In agentic contexts, this process transcends mere static response generation, necessitating dynamic interactions with complex environments to facilitate adaptive decision-making.
  • Preprocessing Pipeline: Post-actor inference, the framework undertakes a sophisticated derivation of auxiliary information critical for RL training. This includes computing log probabilities from the reference policy to quantify behavioral consistency, estimating value functions via the critic model to assess state desirability, and calculating advantage estimates to prioritize high-impact learning signals.
  • Optimization Stage: With the curated dataset from the generative and preprocessing stages—encompassing actor outputs, critic evaluations, and policy gradients—the system initiates iterative parameter updates. This stage harmonizes actor-critic dynamics to refine decision-making strategies, ensuring alignment with task objectives and environmental feedback.
This structured workflow underscores the synergy between generative modeling, value estimation, and policy optimization, cementing RLHF as a cornerstone for developing robust, adaptive intelligent systems.
In the Generative Rollout Phase, contemporary inference engines such as vLLM [412] and SGLang [413] are leveraged to facilitate high-throughput deployment, ensuring efficient realization of generative workflows.
For the Optimization Stage, diverse parallelism paradigms—including data parallelism, pipeline parallelism, and tensor parallelism—are systematically implemented. This is achieved through state-of-the-art distributed training frameworks like Megatron-LM [414] and MegaScale [415], which incorporate sophisticated 3D parallelism. These architectural designs enable seamless orchestration of computational resources, enhancing both scalability and training efficiency across distributed systems.
5.1.1 Taxonomy of RL Infrastructure
While distinct RLHF frameworks broadly adhere to an analogous workflow, notable discrepancies emerge in their architectural configurations. The most pronounced differentiator resides in whether they employ a Single-Controller or Multi-Controller paradigm, a distinction that fundamentally shapes their operational dynamics and scalability profiles.
Within the Single-Controller paradigm, a centralized command hub governs the holistic execution trajectory of distributed programs, enabling users to craft core dataflow functionalities as an integrated workflow. However, the propagation of coordination messages from this central hub to all worker nodes engenders significant dispatch overhead during the execution of large-scale dataflow graphs across sprawling clusters. VeRL and ReaLHF exemplify this architectural model with notable prominence, illustrating both its cohesive design advantages and scalability challenges in high-scale distributed environments.
In the domain of Multi-Controller architectures, each device—hereafter referred to as a worker—is equipped with its own bespoke controller. Cutting-edge distributed systems for LLM training and inference demonstrate a pronounced preference for the multi-controller paradigm. This inclination is primarily attributable to its inherent scalability and the minimal scheduling overhead it incurs.
In the context of implementing the RLHF workflow within a multi-controller architectural framework, practitioners are required to meticulously choreograph the integration of codebase governing inter-device collective communication protocols, computational logic frameworks, and point-to-point data transfer mechanisms across each device’s operational runtime environment. This imperative orchestration gives rise to profoundly nested code architectures, wherein computational routines and data transmission operations become inextricably interwoven. The resultant architectural intricacy poses formidable challenges for development, maintenance, and performance optimization—given that the inter-dependencies between functional components necessitate elaborate debugging methodologies and precision tuning to ensure seamless cross-component synchronization and optimal resource allocation. OpenRLHF stands as a paradigmatic exemplar of such systems, embodying the complex interplay between modular device autonomy and centralized coordination required to sustain efficient RLHF pipeline execution.
5.1.2 Prospective Challenges of RL Infrastructure
  1. 1.
    Long-Text-Generation. The prioritization of long-text generation in reasoning models presents substantial challenges for the generation stage. During Long-CoT generation, the substantial variability in response lengths across different prompts leads to significant GPU under-utilization due to prolonged idle periods. In the context of agentic reinforcement learning (Agentic-RL), the generation stage must not only efficiently produce lengthy responses but also interact dynamically with environments. As environments scale in complexity and diversity, this interaction imposes escalating challenges on generative efficiency and adaptability. For long-text generation, numerous studies have introduced sophisticated rollout strategies to address these issues. For example, Kimi-1.5 [255] employs a partial-rollout framework, eschewing the need for monolithic response generation. Instead, it processes and stores textual segments incrementally, enabling the creation of significantly longer outputs while maintaining rapid iteration cycles. During training, selective exclusion of non-critical segments from loss computation optimizes the learning process, enhancing both operational efficiency and system scalability. Concurrently, Seed-Thinking-1.5 [416] introduces a Streaming Rollout System that decouples model evolution from runtime execution, facilitating dynamic adjustment of on/off-policy sampling. This architectural innovation ensures that generative processes remain responsive to environmental feedback without compromising computational efficiency, thus advancing the state-of-the-art in scalable long-text generation frameworks.
  2. 2.
    Scalable Design for Agent-Environment Interaction. In the realm of agent-environment interaction scenarios [417], the dynamic interplay between agents and their surroundings poses a formidable challenge, primarily stemming from the heterogeneous latencies inherent across diverse environments [418, 408]. To address this, we posit that leveraging an asynchronous message-queue architecture such as the producer-consumer paradigm becomes imperative for effectively managing the temporal discrepancies in agent-environment exchanges. This approach mitigates the adverse effects of variable delays, ensuring robust coordination even in systems characterized by unpredictable response times. Regarding environmental stability, our analysis, rooted in scaling-law principles, underscores the critical role of scalable environmental design in advancing LLM/MLLM-based agentic applications (e.g., DeepResearch [419], DeepSearch [420], and autonomous tooling frameworks like OpenManus [421]). Unlike traditional LLM/MLLM training pipelines, where data resides statically within file systems, agentic environments typically consist of distributed assemblages of HTTP servers. This architectural distinction introduces a nontrivial stability challenge for large-scale Agentic-RL training, as the reliability of such systems hinges on the seamless integration and resilience of networked components. Scaling environments to accommodate growing computational demands while maintaining operational consistency thus emerges as a foundational requirement for realizing the full potential of next-generation agentic systems.
  3. 3.
    Protocol Design and Security Considerations. Developing interfaces between disparate HTTP servers and LLMs can impose substantial overhead, as each integration often requires bespoke engineering efforts. A standardized protocol—such as the Model-Computer Protocol (MCP) [422]—facilitates seamless interaction between LLMs/MLMs and tools without requiring intimate knowledge of tool-specific implementations. This framework streamlines development by mitigating the need for redundant interface design, thereby enhancing interoperability and reducing engineering latency. However, this architectural elegance introduces a critical security imperative: third-party MCP servers may harbor malicious code that could compromise the integrity of the training ecosystem. Such risks necessitate robust safeguards to prevent unauthorized access, code injection, or systemic sabotage, ensuring that the benefits of protocol standardization are not overshadowed by potential vulnerabilities. Proactive measures—including rigorous code audits [423], secure authentication mechanisms [424], and runtime monitoring [425] are essential to maintain the reliability and safety of the integrated system, striking a balance between operational efficiency and defensive resilience.
5.2 Trustworthiness of Reasoning LLMs
As RLLMs grow more capable, their trustworthiness becomes increasingly critical—particularly in domains requiring transparency, factual grounding, and safety. Unlike general LLMs, RLLMs introduce unique trust challenges due to their explicit multi-step reasoning, reliance on external tools, and complex decision traces. Recent work has begun addressing these issues by enhancing interpretability, managing knowledge conflicts, improving robustness, grounding reasoning in evidence, and enabling controllability.
  1. 1.
    Interpretability of Reasoning Processes: RLLMs often produce complex reasoning traces, making it hard to assess whether outputs reflect genuine internal decision-making. Faithfulness metrics have been introduced to evaluate the alignment between explanations and model behavior [426]. Mechanistic analyses explore how reasoning steps are encoded in neurons or attention heads[427, 428]. Other studies examine how gradient dynamics differ when training models for slow, deliberate reasoning [429]. These efforts enhance transparency while preserving performance.
  2. 2.
    Knowledge Conflicts Knowledge inconsistencies arise when internal model memory conflicts with external context, undermining reasoning. Recent surveys categorize such conflicts and highlight their impact on reasoning accuracy [430, 431]. Methods have been proposed to detect and resolve these conflicts using special prompt [432], decoding [433, 434] and alignment [435, 436] technologies, while advanced editing approaches [437, 433, 438] allow updated facts to be incorporated into reasoning traces without full model retraining.
  3. 3.
    Safety & Robustness: Reasoning enables models to reflect on safety but also opens new vulnerabilities. Alignment procedures often reduce reasoning performance, a trade-off termed "safety tax" [439]. Chain-of-thought attacks, such as prompt hijacking and input slowdown, expose structural vulnerabilities in reasoning flows [440, 441]. Backdoor-based jailbreaks further demonstrate that specific reasoning patterns can trigger harmful behavior [442, 443]. To counter this, researchers propose deliberative safety alignment [444], reasoning-based safeguard modules [445], and dedicated safe-CoT benchmarks [446].
  4. 4.
    External Grounding and Truthfulness: RLLMs benefit from grounding reasoning in external evidence to mitigate hallucination. Retrieval-augmented frameworks dynamically adapt context during reasoning [447, 448]. Other prompting techniques first extract chains of evidence from the input before generating an answer [449]. Reinforcement methods reward reasoning paths that include faithful, verifiable external content [450]. Symbolic-verification modules also check reasoning consistency against retrieved data [173].
  5. 5.
    Consistency and Controllability: To improve reliability, RLLMs must generate consistent reasoning across runs and support external control. Dual-model self-reflection and critique techniques help detect and revise flawed reasoning [451]. Prompting strategies that separate high-level plans from step-wise solutions improve coherence [452, 453]. Studies also show that excessively long reasoning chains reduce accuracy, prompting step-length control mechanisms [454].
Across interpretability, safety, grounding, and control, recent research demonstrates growing attention to the trustworthiness of RLLMs. Still, many challenges remain, particularly in balancing reasoning ability with safety, ensuring consistency under open-domain inputs, and aligning long reasoning chains with verified knowledge. Future efforts must continue integrating reasoning supervision, verification, and robust alignment into unified RLLM frameworks.
5.3 Reasoning LLM as Agent
LLMs are increasingly employed as autonomous agents capable of complex reasoning and interaction with external environments. This involves enabling them to actively seek information, utilize diverse tools, and dynamically refine their reasoning. Recent advancements span enhanced search augmentation, broader tool integration, novel training methodologies, and evolving evaluation frameworks.
5.3.1 Search-Augmented Reasoning
Rather than relying solely on internal knowledge, these models are being trained to interact with external search engines to access up-to-date information. Reinforcement learning emerged as a prominent technique for training LLMs to autonomously generate relevant search queries as an integral part of their reasoning chains [455],[456], [457]. Frameworks were developed to incentivize models to invoke search proactively during reasoning, optimizing the interaction based on outcomes rather than predefined processes or distillation [458]. Furthermore, research explored the synergistic potential of combining open-source reasoning LLMs with specialized search tools, demonstrating that such combinations can achieve performance competitive with larger, proprietary systems [459]. The complexity of information retrieval was also addressed through collaborative approaches, where multiple LLM agents work together, potentially using sophisticated search strategies like MCTS, to tackle complex search-based reasoning tasks [460].
5.3.2 Tool Integration Beyond Search
Beyond web search, the scope of tool integration for LLM agents expanded considerably. Efforts focused on enabling LLMs to leverage a wider array of external tools, such as code interpreters and specialized solvers. Self-learning frameworks were introduced, allowing models to acquire tool-using skills through methods like hint-based learning and rejection sampling fine-tuning, effectively teaching them when and how to employ external functionalities [461]. This capability proved valuable in specialized domains; for instance, systems were developed to automate the complex process of modeling and solving Operations Research (OR) problems by translating natural language descriptions into formal mathematical models and generating executable code for solvers [462]. More comprehensive agentic reasoning frameworks emerged, integrating multiple capabilities—including web search [463], code execution [464], and structured reasoning-context memory [465, 466] (e.g., using a mind map structure)—to support more robust and multifaceted problem-solving.
5.3.3 Training Methodologies
Innovations in training methodologies were crucial for developing these sophisticated agentic capabilities. Research demonstrated the effectiveness of reinforcement learning techniques for enhancing reasoning abilities, even in the context of smaller LLMs, suggesting that advanced reasoning is achievable without massive parameter counts [467]. Further emphasizing self-improvement through reinforcement learning, some approaches employ a two-stage paradigm involving initial format tuning to instill structured reasoning patterns incorporating meta-actions, followed by large-scale RL optimization enabling autoregressive search and self-correction with minimal initial supervision [468]. RL-based training was successfully scaled to complex, specialized domains like software engineering, utilizing lightweight, rule-based reward signals to guide the learning process effectively [469]. Advanced RL strategies, such as explicit policy optimization over multiple interaction turns, were explored to improve strategic reasoning performance without requiring initial supervised fine-tuning phases [470]. Additionally, novel training paradigms were developed to cultivate complex reasoning patterns, such as long chains of thought. Bootstrapping methods, for example, enabled the development of these capabilities without relying heavily on advanced teacher models or extensive human annotations, offering a more scalable approach to training deep reasoning [471].
5.3.4 Evaluation Benchmarks and Metrics
As LLM agent capabilities advanced, evaluation methodologies evolved in tandem. New benchmarks were introduced specifically to assess the complex, multi-step reasoning and tool-use abilities required in real-world applications [472]. Interactive platforms and "arenas" gained prominence, providing environments for systematically comparing the performance of different LLM agents on user-defined tasks, often incorporating community voting and feedback mechanisms for a more holistic evaluation [473], [417], [474], [475]. Alongside these new developments, established benchmarks measuring core reasoning abilities (e.g., AIME [295], GPQA [313]) and code generation/execution (e.g., LiveCodeBench [311], SWE-Bench [310]) continued to serve as important yardsticks for tracking progress in foundational capabilities.
5.4 Efficient Reasoning LLM
Despite LongCoT demonstrating strong generalization and reasoning abilities, the autoregressive paradigm of language models imposes a significant reasoning burden, limiting the application of these models in agent-based or edge scenarios [476]. Additionally, numerous studies have shown that language models often exhibit excessive reasoning, with a lot of redundant inference in the thought chains. As a result, various technical approaches have emerged to improve the reasoning efficiency of Reason LLMs. We categorize these approaches into the following categories based on their strategies:
  1. 1.
    Building reasoning-budget-sensitive LLMs: CoD [477] and TALE-EP [478] attempt to impose reasoning budget constraints on Reason LLMs to control the overall reasoning cost. Models like L1 [479], TOPS [480], O1-Pruner [292], and Kimi-k1.5 [255] add length penalties.
  2. 2.
    Building diverse reasoning-length data for post-training: CoT-Valve [481] synthesizes diverse-length thought chain data through data interpolation. TOPS samples corresponding versions using a budget-sensitive data model, and C3oT [482] compresses the original LLM output to obtain shorter Short-CoT and trains them jointly.
  3. 3.
    Using external models or switching mechanisms [483] to intervene in the reasoning scope: Routellm [484] introduces multiple routers to find the most suitable reasoning model for a given problem, while Self-REF [485] uses confidence to route the model’s reasoning difficulty. Methods like JudgeLRM [486] and DeepSeek-GRM [166] attempt to construct generalized Reward Models for efficient test-time scaling, eliminating irrelevant reasoning paths.
  4. 4.
    Using efficient representations to execute reasoning: TokenSkip selects data based on the importance of tokens and executes reasoning using compressed, more concise thought chain representations. COCONUT [487] attempts to execute more efficient reasoning in the latent space. ICoT-KD [488] and CCoT [488] try to build more efficient reasoning strategies in the hidden space, while Token Assorted combines hidden space reasoning with text-based reasoning to balance interpretability and efficiency. Heima aims to build hidden-space reasoning for multimodal models.
  5. 5.
    Non-autoregressive reasoning models: Diffusion-LM avoids a large number of autoregressive prediction steps due to its non-autoregressive nature, enabling efficient reasoning. LLaDA [489] constructs a diffusion-based LM with an enormous parameter size, achieving scalability based on a non-autoregressive language model architecture. Diffusion-of-Thoughts [490] uses post-training strategies to transform an LLM into a denoising process for token prediction, significantly improving reasoning efficiency.
  6. 6.
    New architectures for efficient reasoning: One of the key bottlenecks in long thought-chain prediction is the computational cost of ultra-long context. Mamba [491] and RWKV [492], among others, use linear attention or state space techniques to effectively enhance the model’s reasoning efficiency.